# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

### v3

@@ Abstract.
The potential impact of an academic paper is determined by various factors, including its popularity and contribution. Existing models usually estimate original citation counts based on static graphs and fail to differentiate values from nuanced perspectives. In this study, we propose a novel graph neural network to Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution values (called DPPDCC). Given a target paper, DPPDCC encodes temporal and structural features within the constructed dynamic heterogeneous graph. Particularly, to capture the knowledge flow, we emphasize the importance of comparative and co-cited/citing information between papers and aggregate snapshots evolutionarily. To unravel popularity, we contrast augmented graphs to extract the essence of diffusion and predict the accumulated citation binning to model conformity. We further apply orthogonal constraints to encourage distinct modeling of each perspective and preserve the inherent value of contribution. To evaluate models’ generalization for papers published at various times, we reformulate the problem by partitioning data based on specific time points to mirror real-world conditions. Extensive experimental results on three datasets demonstrate that DPPDCC significantly outperforms baselines for previously, freshly, and immediately published papers. Further analyses confirm its robust capabilities. We will make our datasets and codes publicly available.
Impact Prediction, Hierarchical and Heterogeneous Graph, Graph Neural Network, Disentangled Representation Learning, Contrastive Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Theory of computation Dynamic graph algorithmsccs: Mathematics of computing Graph algorithms
@@ 1. Introduction
In paper retrieval and recommendation, ranking papers based on their scientific impact can aid researchers to delve into intricate research efforts. This is particularly significant considering the exponential annual growth in the number of published papers (Lo et al., 2020; Chu and Evans, 2021; Xue et al., 2023). Given the inherent difficulties in quantifying the precise numerical value of scientific impact, citation count is regularly employed as a rough approximation (Evans and Reimer, 2009; Sinatra et al., 2016; Jiang et al., 2021). Nevertheless, the current ranking depends on the cumulative number of citations received to date, which merely represents the impact within the prior research environment. Therefore, it becomes imperative to cultivate a sophisticated model that can assess the potential impact of a research paper, placing significant emphasis on the future (Wang et al., 2013).
@@
At a given point in time (e.g., at present), estimating the anticipated increment in citations can effectively facilitate a fair comparison of the potential impact between long-published papers and newly published ones. Actually, the citations received by papers are not solely derived from the influence of their contributions, but are also affected by a multitude of other factors (Bornmann and Daniel, 2008; Case and Higgins, 2000). Hence, rather than directly forecasting original citation counts, it is more effective to disentangle citation increments received from extrinsic factors and the actual contributing factor, making impact prediction more practical and interpretable. Afterwards, we can identify exceptional works amidst the vast array of publications by discerning the potential contributing impact. In this work, we emphasize the popularity factor of a paper as a prominent extrinsic factor affecting the increase of citations. More precisely, when referring to the popularity factor of a paper, we are considering its citation diffusion patterns and the collective conformity degree within the citation network.
@@
Regarding citation diffusion, we posit that influential papers assume an amplifying effect in shaping the process of information dissemination within the citation network (Page et al., 1998; Wu et al., 2019). This leads to increased visibility of papers connected to them. In other words, papers cited by highly-cited papers will receive significantly more citations than others. The statistics of the dataset provides empirical support. For instance, in the field of Computer Science, the former papers will receive an average of 1.5475 citations (after applying a logarithmic transformation), while the latter ones will only receive an average of 0.8580 citations. Towards collective conformity, existing studies have found that some researchers may tend to cite papers associated with well-known entities, irrespective of their actual relevance or genuine contribution (Thornley et al., 2015). This phenomenon is substantiated by the future disparity (2.1917 vs. 0.8009) in papers with different accumulated citations, which aligns with the concept of Matthew effect in the academic domain (Allison and Stewart, 1974).
@@
Therefore, by separating out the influence of the popularity, i.e., citation diffusion and collective conformity, we can gain a closer approximation to the genuine contributing impacts of papers. In essence, we aim to disentangle the citation increment into diffusion, conformity, and contribution values (as shown in Figure 1 (a)).
@@
Figure 1. (a) We disentangle the citation increment into diffusion, conformity, and contribution values for better interpretability. (b) To imitate the evaluation of future dynamics of the citation context, we split the dataset by the observation time point and categorize the samples into previous, fresh, and immediate papers in terms of their publication time. (c) We represent the heterogeneous subgraph schema with all metadata, where each type of relation is bidirectional.
@@
Typically, conventional research (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022) on citation prediction imposes stringent constraints to select favorable samples by removing lowly-cited papers. Moreover, the division of these data into training, validation, and testing is carried out randomly, disregarding the chronological order of paper publications (Ruan et al., 2020; Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023). However, this problem setting deviates from real-world scenarios, possibly causing a substantial distribution shift. Hence, the model may face challenges in handling recently published papers. Random data partitioning disregards new citation network updates, hampering the model’s practicality and robustness in upcoming real-world test. This limitation impedes the exploration of the potential value of papers reactivated in the evolved context (Van Raan, 2004).
@@
In contrast, our work emulates reality by performing a novel problem setting. We include all papers without strict filtering to maintain consistency with the data distribution of real-world citation network. As illustrated in Figure 1 (b), we partition the dataset based on particular observation time points. This can introduce fresh papers that are not part of the training set (previous papers) into the test set. We further consider immediate papers (published at the test time point) that closely align with cold-start scenarios to validate the robustness of our model. Thus, papers in the test set can be divided into previous, fresh, and immediate papers, facilitating a more practical evaluation. Notably, previous papers have different contexts and future increments at the train and test time points. By employing this approach, our model is incentivized to fit the dynamic citation context, thereby enhancing its generalization.
@@
Existing graph disentanglement studies usually aim at decomposing abstractive hidden factors from either static or homogeneous graphs (see Section 2) (Ma et al., 2019a; Yang et al., 2020; Wang et al., 2020). Instead, our focus lies in disentangling the categorical characteristics of citation behaviors within dynamic and heterogeneous citation graphs. In this study, tailored to each target paper, we construct a Dynamic Heterogeneous Graph from the citation network. The graph consists of a sequence of heterogeneous snapshots (see Figure 1 (c)). This graph retains both the content of the paper and its dynamic citation context. Based on the proposed graph, we devise a model that could Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution Values (called DPPDCC). This model utilizes Citation-aware Relational Graph Convolutional Networks (R-GCNs) to capture the structural features of the input sequence of snapshots. The outputs are then passed through a Transformer Encoder, which allows for the incorporation of temporal information. Specially, we design a novel module in modified R-GCN to integrate the comparative and co-cited/citing information between paper nodes. Furthermore, we propose a Type-specific Attention Readout mechanism to aggregate information from each snapshot effectively. Upon acquiring the final representation of the target paper, we proceed to disentangle the predicted citation increment into diffusion, conformity, and contribution values individually. For the diffusion, we employ Triplet Citation-aware Graph Contrastive Learning to capture the popularity influence inherent in the information diffusion process within the citation network. For the conformity, we introduce an auxiliary task to classify bins of accumulated citations. Additionally, we impose an orthogonal constraint on the representations of disentangled perspectives in pair, guaranteeing that they encode distinct features without overlap. Extensive experiments are conducted on three fields of the real-world scientific dataset S2AG (Kinney et al., 2023). The results illustrate the superior performance of DPPDCC compared to baseline models across previous, fresh, and immediate papers. Ablation tests further highlight the significance of the proposed components. Additionally, visualizations offer insights into the model’s interpretability.
@@
$\bullet$ We revise the problem formulation of impact prediction to better align with real-world scenarios. Without filtering the predicted samples by their citation counts, we partition the datasets based on specific time points to maintain the dynamics of the citation context. Moreover, we employ disentangled representation learning to extract the contributing impacts of papers.
$\bullet$ We propose a novel model called DPPDCC that first encodes comparative and co-cited/citing information evolutionarily within the Dynamic Heterogeneous Graph through Citation-aware Graph Neural Network Encoder for potential impact prediction. It further disentangles the citation increment into diffusion, conformity, and contribution values.
$\bullet$ Experimental results show significantly superior to existing baselines for previous, fresh, and immediate papers. Further analyses illustrate that our model can reasonably make predictions.
@@ 2. Related Work
We review three lines of related work: Citation/Cascade Prediction, Dynamic and Heterogeneous Graph Neural Network (GNN), and Disentangled Representation Learning.
2.1. Citation/Cascade prediction
Citation prediction is a vital sub-task of automated academic evaluation as it enables the estimation of potential impact. It is closely examined within cascade prediction tasks, which share analogous graph structures and objectives (Zhou et al., 2021). The primary goal of cascade prediction is to anticipate the popularity of a post based on the interactions between users. Similarly, in the context of citation networks, the paper can serve as the post, while the authors can serve as the users. Citation count prediction places a focus on the content and distinctive features of individual papers, heavily relying on comprehensive feature extraction techniques. In contrast, cascade prediction delves into the intricate relationships between various entities, including papers and authors, directly exploiting the information of graph structure.
@@ 2.1. Citation/Cascade prediction
Their approaches can be broadly classified into three categories: stochastic models, feature-based models, and deep learning models. Stochastic models can predict future citation counts by fitting the curve of citation counts (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997). Recently, machine learning models have shown promising results by utilizing manually extracted features from various meta-data (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Guille et al., 2013; Tatar et al., 2014). More presently, deep neural networks have dominated this task by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023). For cascade prediction, graph embeddings, sequence models, and GNN models are employed to extract structural information from the underlying graph, and then temporal information is encoded with sequence models (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022).
However, existing models for citation and cascade prediction exhibit sub-optimal performance. They fail to fully exploit the valuable information presented in paper content and scientific context within the citation network simultaneously. Furthermore, their strict data selection and random data splitting strategies may hinder their practicality. Many valuable samples will be filtered and the temporal information within dynamic context will be neglected. Thus, models in this setting are more prone to face distribution shift issues. In contrast, our study proposes to specifically revise these settings and utilize both content and dynamic context for impact estimation. We also emphasize retaining genuine contributions to bolster the robustness of our models.
2.2. Dynamic/Heterogeneous Graph Neural Networks
GNNs (Sperduti and Starita, 1997; Gori et al., 2005; Scarselli et al., 2008; Wu et al., 2020) are widely applied to handle non-Euclidean data like graphs. Recently, research interest has surged in dynamic graphs, which incorporate temporal information, as well as in heterogeneous graphs, which involve multiple node types or edge types. These types of graphs are more prevalent in practical scenarios and encapsulate more intricate information, compared to static homogeneous graphs, driving the formulation and refinement of diverse methodologies within the academic realm.
Citation network can be regarded as dynamic and heterogeneous graphs, as it evolves over time and encompasses diverse entities such as papers, authors, and venues. Addressing the complexities within this network involves two primary challenges: (1) How to encode the temporal target-centric information within a single snapshot? (2) How to model distinct characteristics of information diffusion in citation networks? To overcome these challenges, our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers. These graphs are constructed from the citation network on an annual basis and incorporate multiple metadata nodes beyond papers themselves. (1) To encode the temporal target-centric information within a single snapshot, we employ a type-aware attention readout. It aggregates snapshot information considering paper type and publication time. (2) To model distinct characteristics of information diffusion in citation networks, we propose a novel GCN module CompGAT. This module focuses on capturing comparisons and co-cited/citing information between papers, drawing inspiration from bibliometrics theories (Wu et al., 2019).
2.3. Disentangled Representation Learning
Recently, there has been a growing interest in applying disentangled representation learning to graph learning tasks. Given that graph data encompasses both structural and attributed information, numerous new methodologies have emerged to address this complexity. In the feature space, inspired by capsule networks, approaches have emerged to partition node features into multiple hidden channels (Ma et al., 2019a). Some methods introduce additional constraints to enable effective separation (Liu et al., 2020). Moving into the structural space, subsequent models strive to factorize the input graphs into multiple subgraphs, to facilitate distinct message-passing strategies (Yang et al., 2020; Li et al., 2021). Recent models have expanded their scope to isolate causal and biased information presented in both the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022). These models commonly employ techniques to generate masks within the adjacency matrix and node features. This process aims to identify the crucial elements of the graph within the structural and attributed spaces. In addition, certain studies extend disentangled GNNs to specialized scenarios such as recommendation systems, handling intricate graph structures in heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022). Besides the masking strategy considering solely the graph structures and node features, they integrate classical methods like VAE and adversarial attack learning from CV and NLP to handle task-specific information from a broader perspective.
In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by factors related to its popularity. These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect through collective conformity. Based on the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, seeking a more comprehensive estimation. Specifically, we devise dedicated auxiliary tasks aimed at extracting values associated with these respective factors.

### v4

# Predicting Scientific Impact Through Diffusion, Conformity, and Contribution Disentanglement

@@ Abstract.
The scientific impact of academic papers is influenced by intricate factors such as dynamic popularity and inherent contribution. Existing models typically rely on static graphs for citation count estimation, failing to differentiate among its sources. In contrast, we propose distinguishing effects derived from various factors and predicting citation increments as estimated potential impacts within the dynamic context. In this research, we introduce a novel model, DPPDCC, which Disentangles the Potential impacts of Papers into Diffusion, Conformity, and Contribution values. It encodes temporal and structural features within dynamic heterogeneous graphs derived from the citation networks and applies various auxiliary tasks for disentanglement. By emphasizing comparative and co-cited/citing information and aggregating snapshots evolutionarily, DPPDCC captures knowledge flow within the citation network. Afterwards, popularity is outlined by contrasting augmented graphs to extract the essence of citation diffusion and predicting citation accumulation bins for quantitative conformity modeling. Orthogonal constraints ensure distinct modeling of each perspective, preserving the contribution value. To gauge generalization across publication times and replicate the realistic dynamic context, we partition data based on specific time points and retain all samples without strict filtering. Extensive experiments on three datasets validate DPPDCC’s superiority over baselines for papers published previously, freshly, and immediately, with further analyses confirming its robustness. Our codes and supplementary materials can be found at Github.
Impact Prediction, Citation Network, Dynamic Heterogeneous Graph, Graph Neural Network, Contrastive Learning
copyright: acmlicensedjournalyear: 2024conference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679546isbn: 979-8-4007-0436-9/24/10ccs: Information systems Data mining
@@ 1. Introduction
In paper retrieval and recommendation, ranking papers based on their scientific impact can aid researchers in delving into intricate research efforts. This is particularly significant considering the exponential annual growth in the number of published papers (Lo et al., 2020; Chu and Evans, 2021; Xue et al., 2023). Given the inherent difficulties in quantifying the precise numerical value of scientific impact, citation count is regularly employed as a rough approximation (Evans and Reimer, 2009; Sinatra et al., 2016; Jiang et al., 2021; Geng et al., 2022; Yang and Han, 2023). Nevertheless, the current ranking depends on the cumulative number of citations received to date, which merely represents the impact within the prior research environment. Therefore, it becomes imperative to cultivate a sophisticated model that can assess the potential impact of a research paper, placing significant emphasis on the future (Wang et al., 2013).
@@
At a given point in time (e.g., at present), estimating the anticipated increment in citations can effectively serve as a better approximate measure of the potential impact, allowing for a fair comparison between long-published papers and newly-published ones. Actually, the citations received by papers are not solely derived from the influence of their contributions but are also affected by a multitude of other factors (Bornmann and Daniel, 2008; Case and Higgins, 2000). Hence, rather than directly forecasting original citation counts, it is more effective to disentangle citation increments received from extrinsic factors and the actual contributing factor, making impact prediction more practical and interpretable. Afterwards, we can identify exceptional works amidst the vast array of publications by discerning the potential contributing impact. In this work, we emphasize the popularity factor of a paper as a prominent extrinsic factor affecting the increase of citations. More precisely, when referring to the popularity factor of a paper, we are considering its citation diffusion patterns and the collective conformity degree within the citation network.
@@
Regarding citation diffusion, we posit that influential papers assume an amplifying effect in shaping the information dissemination process within the citation network (Page et al., 1998; Wu et al., 2019). This leads to increased visibility of papers connected to them. In other words, papers cited by highly-cited papers will receive significantly more citations than others. The statistics of the datasets provide empirical support. For instance, in the field of Computer Science, the former papers will receive an average of 1.5475 citations (after applying a logarithmic transformation), while the latter ones will only receive an average of 0.8580 citations. Towards collective conformity, existing studies have found that some researchers may tend to cite papers associated with well-known entities, irrespective of their actual relevance or genuine contribution (Thornley et al., 2015). This phenomenon is substantiated by the future disparity (2.1917 vs. 0.8009) in papers with different accumulated citations, which aligns with the concept of the Matthew effect in the academic domain (Allison and Stewart, 1974).
@@
Therefore, by separating out the influence of the popularity, i.e., citation diffusion, and collective conformity, we can better approximate the genuine contributing impacts of papers. In essence, we aim to disentangle the citation increment into diffusion, conformity, and contribution values (as shown in Figure 1a).
@@
Figure 1. (a) We disentangle the citation increment into diffusion, conformity, and contribution values for better interpretability. (b) To imitate the evaluation of future dynamics within the citation context, we split the dataset by the observation time point and categorize samples into previous, fresh, and immediate papers in terms of publication time. (c) We represent the heterogeneous subgraph schema with all metadata, where each relation type is bidirectional.
@@
Typically, conventional research (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022b; Xu et al., 2022; He et al., 2023) on citation prediction imposes strict constraints to select favorable samples by removing lowly-cited papers. Additionally, their division of data into training, validation, and testing sets is typically performed randomly, ignoring the chronological order of paper publications (Ruan et al., 2020; Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023). Moreover, some studies focus solely on the most recently published papers, neglecting the value of those published previously (Jiang et al., 2021; Geng et al., 2022; Yang and Han, 2023). These problem settings deviate from real-world scenarios, potentially causing a substantial distribution shift. Consequently, models may struggle to handle papers published at different times. Overall, the random data partitioning setting disregards new updates to the citation network, hampering the model’s practicality and robustness in real-world applications, while the static problem setting impedes the exploration of the potential value of papers reactivated in the evolving context (Van Raan, 2004).
@@
In contrast, our work emulates reality by revising the dynamic problem setting of previous studies (He et al., 2023). We include all papers without strict filtering to maintain consistency with the real-world distribution of citation network data. As illustrated in Figure 1b, we partition the dataset based on specific observation time points, ensuring distinct training and testing windows. This methodology introduces fresh papers—excluded from the training set (previous papers) —as part of the test set, serving both as samples and context. We further consider immediate papers (published at the test time point) that closely align with cold-start scenarios, facilitating a more practical evaluation. Notably, previous papers have different contexts and future increments at the train and test time points. Adopting this strategy incentivizes our model to fit the dynamic citation context, thereby enhancing its generalization.
@@
Existing graph disentanglement studies typically aim at implicitly decomposing abstractive hidden factors from either static or homogeneous graphs (see Section 2) (Ma et al., 2019a; Yang et al., 2020; Wang et al., 2020). In contrast, our focus is explicitly disentangling citation behaviors’ categorical characteristics within dynamic and heterogeneous citation graphs. For each target paper, we construct a Dynamic Heterogeneous Graph from the citation network, consisting of a sequence of heterogeneous snapshots (see Figure 1c), which retains both the content of the paper and its dynamic citation context. Based on this graph, we devise a model named DPPDCC that Disentangles the Potential impacts of Papers into Diffusion, Conformity, and Contribution values. The model first employs a Citation-aware Graph Neural Network Encoder (CGE) to extract distinctive structural and temporal information alternately. The Popularity-aware Disentanglement Module (PDM) then separates the values derived from different factors through three simultaneous auxiliary tasks. Specifically, the model utilizes Citation-aware Relational Graph Convolutional Networks (R-GCNs) to capture the structural features of the input snapshot sequence and a Transformer Encoder to incorporate temporal information. A novel message-passing module within the R-GCN, CompGAT, is designed to integrate comparative and co-cited/citing information between papers. Additionally, the Type-specific Attention Snapshot Readout is proposed to aggregate information from each snapshot evolutionarily. Upon acquiring the final representation of the target paper, we disentangle the predicted increment into diffusion, conformity, and contribution values. For diffusion, we employ Triplet Citation-aware Graph Contrastive Learning to capture the popularity influence inherent in the information diffusion process within the citation network. For conformity, we introduce an auxiliary task to classify bins of accumulated citations. Furthermore, orthogonal constraints are imposed on the representations of disentangled perspectives in pairs, ensuring they encode distinct features without overlap. Extensive experiments conducted on three fields of the real-world scientific dataset S2AG (Kinney et al., 2023) demonstrate DPPDCC’s superior performance compared to baseline models across previous, fresh, and immediate papers. Ablation tests further highlight the significance of the proposed components, and visualizations provide insights into the model’s interpretability.
@@
$\bullet$ We revise the problem formulation of impact prediction to better align with real-world scenarios. Without filtering predicted samples by their citation counts, we partition the datasets based on specific time points to maintain the dynamics of the citation context. Moreover, we employ disentangled representation learning to extract the contributing impacts of papers.
$\bullet$ We propose a novel model, DPPDCC, which first encodes comparative and co-cited/citing information evolutionarily within the Dynamic Heterogeneous Graph through a Citation-aware Graph Neural Network Encoder. It further disentangles the citation increment into diffusion, conformity, and contribution values.
$\bullet$ Experimental results show that DPPDCC notably outperforms existing baselines for previous, fresh, and immediate papers, and further analyses confirm our model’s reasonable prediction ability.
@@ 2. Related Work
We review three lines of related work: Citation/Cascade Prediction, Dynamic and Heterogeneous Graph Neural Networks (GNNs), and Disentangled Representation Learning.
@@
Citation prediction is a vital sub-task of automatic academic evaluation as it enables the estimation of scientific impact. It is closely examined within cascade prediction, which shares analogous graph structures and objectives (Zhou et al., 2021). Their approaches can be classified into three categories: stochastic, feature-based, and deep learning models. Stochastic models predict future citation counts by fitting citation curves (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997). Recently, machine learning models have demonstrated promising results using manually extracted features from various metadata (e.g. content, author, venue) (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Dong et al., 2015). More presently, deep neural networks have dominated by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Cohan et al., 2020; Xue et al., 2023). For cascade prediction, graph embeddings, sequence models, and GNN models extract structural information from the graph, with temporal information encoded by sequence models (Li et al., 2017; Cao et al., 2017; Xu et al., 2022; Cheng et al., 2023). However, most existing models exhibit suboptimal performance, failing to simultaneously exploit the valuable information presented in paper content and the scientific context within the citation network. Furthermore, their strict data selection and static random data splitting strategies may hinder their practicality. In contrast, our study revises the problem setting of impact prediction in the dynamic context (He et al., 2023; Yan et al., 2024). We ensure non-overlapping training and testing observation windows, retaining the real data distribution without sample filtering. We also utilize both paper content and dynamic context, and further introduce disentangled representation learning to provide a more practical impact estimation.
GNNs (Sperduti and Starita, 1997; Gori et al., 2005; Scarselli et al., 2008; Wu et al., 2020) are widely applied to handle non-Euclidean data like graphs. Dynamic graphs incorporate temporal information, while heterogeneous graphs involve multiple types of nodes or edges. Dynamic graphs can be divided into snapshots at different time points. Previous models like DGCN (Manessi et al., 2020), Dysat (Sankar et al., 2020), and ROLAND (You et al., 2022) typically encode the snapshots with static GCN(s) as the structural encoder, followed by sequential models like RNN or Transformer Encoder. In heterogeneous graphs, R-GCN serves as a prominent model (Schlichtkrull et al., 2018). It conducts independent message passing within different relations and aggregates them to update node representations. Our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers constructed from the citation network. It utilizes modified R-GCNs as structural encoders and a multi-layer transformer as the temporal encoder. Additionally, we incorporate bibliometric theory to model distinct features within the citation network.
Disentangled representation learning aims to identify and disentangle the underlying explanatory factors (Bengio et al., 2013). The existing efforts primarily focus on CV and NLP, utilizing Variational Auto-encoders (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021). In graph learning tasks inspired by capsule networks, they partition node features into multiple hidden channels for implicit disentangled modeling (Ma et al., 2019a). Subsequent models strive to factorize the input graphs to facilitate distinct message passing (Yang et al., 2020). Recent models expand their scope to isolate causal and biased information presented in the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022). In addition, certain studies extend the disentangled GNNs to specialized scenarios like recommendations and healthcare, addressing more complex graph structures involving either heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022; Zhang et al., 2023b, a). In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by popularity factors. These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect as collective conformity. Building upon previous disentangled studies in recommendations (Zheng et al., 2021; Chen et al., 2022a), we regard the final predicted values as mixed results of multiple factors and introduce auxiliary tasks for separate encoding of different perspectives. Leveraging the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, aiming for a more rational estimation through explicit modeling.